uestion 1) 2 points .Suppose a person wants to allocate income () between two products H (on the horizontal axis) and S (on the vertical axis) are given: PH-1:Ps-8:1-100: a. Write the equation of the budget line b. Compute the horizontal intercept& write the interpretation of it c. Compute the vertical intercept& write the interpretation of it d. Compute the slope and write the interpretation of it Question 2) 2 points .Consider information in question 1 to be the baseline scenario. Now we want to create two changes and compare them to the baseline scenario Scenario 1 Government imposes a $0.50 tax on the per unit price of product H Scenario 2 The producers give a discount of SO.30 to consumers if H is purchased more than 60 units a. Find the equation of the budget line, slope, horizontal intercept, vertical intercept for b. scenario 1 Find the equation of the budget line, slope of Budget Line scenario 2 Draw two graphs: Graph 2.1 compares the baseline budget constraint to scenario 1 and Graph 2.2 compares the baseline budget constraint to scenario 2 c.
Answer #1

Q1) a) Budget line represents different combinations of goods that consumer can purchase with his given money income and given prices of goods. The equation of the budget line in this case is given as follows:

(H X PH) + (S X PS) = I

where, H is the quantity of good H.

PH is the price of good H which is given as 1 in this case.

S is the quantity of good S.

PS is the price of good S which is given as 8 in this case.

I is the income which is given to be 100 in this case.

Putting the value of prices in the above equation,

(H X 1) + ( S X 8) = 100

H + 8S = 100 is the equation of the budget line

b) The horizontal intercept can be calculated by considering the case when S = 0. If S = 0, then H = 100. This means that if consumer spends his entire income on good H, he can get 100 units of good H and 0 units of good S. This is the interpretation of the horizontal intercept that if consumer spends nothing on good S and spends his entire income of good H, he can consume 100 units of good H.

When S = 0, H = 100 (by putting S = 0 in the budget constraint)

c) The vertical intercept can be calculated by considering the case when H = 0. When H = 0, the consumer can consume 12.5 units ( 8S = 100, so S = 12.5) of good S. The interpretation of this is as follows, if the consumer spends nothing on H and spends his entire income on consumption on good S, he can consume 12.5 units of good S and 0 units of good H.

When H = 0, 8S = 100 ( by putting H = 0 in the budget constraint)

S = 100/8

S = 12.5

d) The slope of the budget line has an important economic interpretation. It measures the rate at which consumer can exchange one good for the other. It also measures the opportunity cost of getting one good.

The slope of the budget line = Change in good S / Change in good H

The slope of the budget line is also equal to the price ratio of two goods that is,

slope of budget line = Price of good H / Price of good S

= 1 / 8

This means that the consumer can get 8 units of good H by sacrificing or giving up 1 unit of good S.